OpenStudy (anonymous):

integrate e^x / 1+e^x

OpenStudy (babyslapmafro):

use u-substitution u=(1+e^x)

OpenStudy (babyslapmafro):

are you familiar with u-substitution?

OpenStudy (kc_kennylau):

\(u=1+e^x\), \(du=e^xdx\)

OpenStudy (anonymous):

oh! thank you!

OpenStudy (kc_kennylau):

no problem :)

OpenStudy (anonymous):

I got ln 1+e^x+c is this correct?

OpenStudy (kc_kennylau):

nope :/

OpenStudy (kc_kennylau):

Wolfram|Alpha is a great website for checking answer :)

OpenStudy (kc_kennylau):

\[\hspace{11pt}\int\frac{e^x}{1+e^x}dx\]\[=\int\frac1udu\]\[=\ln u+c\]\[=\frac1{1+e^x}+c\]

OpenStudy (isaiah.feynman):